Status Report

NASA Space Station On-Orbit Status 13 Feb 2004

By SpaceRef Editor
February 13, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 13 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

As is standard practice after Russian dockings, FE Kaleri took two photos of the Service Module (SM) aft-end docking assembly (StA) used for the Progress M1-11/13P linkup.  These images will be used to refine current understanding of docking conditions.   [The objective was to take photo imagery of the scratch or scuff mark left by the head of the docking probe on the internal surface of the docking cone ring.  As other cosmonauts before him, Sasha used the Kodak 760 digital still camera to take two pictures with the hatch closed down and downlinked them later (3:45am EST) via OCA.]

In the DC-1 “Pirs” module, Alex Kaleri removed and replaced the PZU/TA765B ROM (read-only memory) unit of the BITS2-12 onboard telemetry system’s LKT local temperature sensor commutator, using a spare delivered on 13P.

CDR Michael Foale completed the microbial (bacterial and fungal) T+5 Day analysis of air and surface samples collected on 2/9 in Lab, Node and SM.   [The T+2 Day analysis had been deleted for SSK and MAS operations.  Bacterial and fungal air samples were taken at two locations in each module.  The colony growths incubated on the SSK (surface sampler kit) sampling slides and the MAS (microbial air sampler) Petri dishes were visually analyzed by comparison to an uplinked colony density chart and photographed; all microbiology sample data were downloaded to the medical equipment computer (MEC), and the equipment was then stowed.  There is some evidence that two Petri dishes (one fungal, one microbial) had been contaminated and, if so, may have contaminated subsequent SM and Node samples via the sampler gear.  The crew was instructed to use a procedure for cleaning the sampler with biocide wipes after completion of the current sampling session and to discard the used slides and Petri dishes.]

On the Elektron O2 generator, Sasha Kaleri removed the faulty differential pressure regulator (RPD) and replaced it with a spare unit.  The electrolysis machine is now operating nominally in 20 amps. mode.   [During the R&R, one of the SM smoke detectors initiated a false fire event, triggering the standard ISS fire isolation response. The offending smoke detector is currently inhibited, and the nominal ECLSS (environmental control & life support system) configuration is restored.]

Kaleri and Foale also completed the long-overdue installation of a coolant line on the SKV-2 air conditioner, delivered on 13P.   [The new line connects the SKV-2’s BTA heat exchanger to an internal thermal control loop.  On 12/17/03 the crew had not been able to connect the old line with the SKV-2’s BTA.  SKV-2 is scheduled to be activated early next week.  SKV-1, currently off, will experience occasional shutdowns when operated in conjunction with internal cooling loop #1 (KOV-1) due to low Khladon (Freon) temperatures.  Because of the raised humidity level (i.e., dew point) in the cabin atmosphere, the U.S. Airlock CCAA (common carbon air assembly) was activated yesterday, to prevent condensate accumulation on the CCAA heat exchanger.]

The two crewmembers took turns in performing the mandatory Russian pre-EVA MedOps procedure MO-6 (hand-cycle ergometry), each providing pomosh’ (assistance) to the other during the individual’s 30-min. test run.   [Previous cosmonauts have shown noticeable decrease in arm muscle tone, and TsUP/IBMP (MCC-Moscow/Institute of Biomedical Problems) physical fitness experts have made the tolerance handgrip test analysis (hand ergometry) a standard pre-Orlan EVA requirement.  MO-6 today required the CDR, later the FE, to don the ECG (electrocardiogram) biomed harness, attach three skin electrodes and plug the harness into the PKO medical exam panel on the cycle ergometer.  The exercise itself started after 10 seconds of complete rest, by manually rotating the cycle’s pedals, set at 150 W, backwards until “complete exhaustion”.]

Mike used the U.S. CDMK (carbon dioxide monitoring kit) to take the weekly reading of the cabin air’s CO2 partial pressure in the SM, for calldown to the ground (along with the battery status) for use in trending analyses, then restowing the CDMK at the SM Central Post.

The crew completed their physical exercise program on CEVIS bike, VELO ergometer and RED exerciser.  Mike then downloaded the accumulated data files from the exercise equipment to the MEC via memory card and RED log entries, for downlink on OCA comm.  The CDR also performed the periodic (weekly) transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, and then deleted them on the HRM.  

Foale and Kaleri reviewed the uplinked preliminary timeline for the EVA-9 on 2/26; they also studied the applicable Russian operations data files (RODFs), supported by tagup with ground specialists via S-band.   [During the spacewalk, the CDR will be EV1, the FE EV2.  Opening of the EV hatch and egress from the DC-1 will be at ~ 4:20pm, right after the first of four eclipses (i.e., orbital night periods, during which Orlan EVAs, as opposed to EMU EVAs, come to a stop for crew rest).  Estimated duration of the EVA is 5 hours 31 minutes.  The very busy timeline lists 46 individual activity steps, including translations to worksites.]

MCC-H uplinked the new EVA-required model updates for the onboard DOUG (Dynamic Operational Ubiquitous Graphics) applications, to run on the new A31p ThinkPad and older 760XD laptops.  The DOUG files now include the pertinent external payloads (Kromka, SKK, MPAC/SEED, etc.)

The rather intricate sleep cycle shifts developed by U.S. and Russian specialists for the upcoming EVA activities will begin tomorrow (2/14), with sleep time starting four hours earlier than normal (12:30pm EST instead of 4:30pm), followed by a four-hour earlier wakeup on 2/15 (Sunday), i.e., to 9:00pm from 1:00am.  This schedule will be retained through the EVA/Soyuz ingress demo on 2/19, then undergo new shifting.

Sasha performed the daily routine maintenance of the SM’s SOZh life support system and prepared the daily IMS (inventory management system) “delta file” for automatic export/import to update the databases, while Mike conducted the regular routine status checkup of autonomous Increment 8 payloads.

Kaleri also held the weekly tagup with IMS specialists at MCC-M, discussing open issues concerning identification of equipment and storage locations for the IMS databases.   [Issues today included calldown of the locations of some specific items and the necessary updating of the IMS database after the recent Matryoshka electronics installation.]

At about 2:45pm, the crew had their weekly teleconference with the ISS Flight Director at MCC-H.

Moscow now concurs with the U.S. procedure to evacuate the Lab window inter-pane volume (“Volume D”).  The procedure will be placed on the crew’s task list tonight and could be performed as early as this weekend.  If not, the activity is penciled into the crew’s schedule for 3/5. 

Today’s CEO (crew earth observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Seoul, South Korea (looking a touch right of nadir for the city center.  Seoul became the capital city in 1394, and with 10.3 million people, has one of the highest densities in the world), Smog, North China plain (Dynamic event.  Looking right onto the North China Plain, one of the largest population centers in the world.  Shooting high obliques for this polluted air mass which has been accumulating), Taiwan Smog (pass over Taiwan: looking right to document smog mass, especially any margins that may have been visible), Bahamas (detailed images of coral reef distribution were requested for the mapping project), Pyrenees snow (Dynamic event.  Clear weather over this snow-covered range), Los Angeles, CA (nadir pass; looking left and right), and Laysan, Hawaiian island chain (a detailed image of the coral reef around this 2 sq-mile atoll was requested.  Due to the island’s remoteness, five endemic species of bird evolved on Laysan).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:26pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (20A).  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S.

  • CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off (repair now completed).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.4; ppO2 (mmHg) — 149.3; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 21.6.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 745.16; temperature (deg C) — 23.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 747.75; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 747.85; temperature (deg C) — 23.9; shell heater temp (deg C) — 23.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.1
  • PMA-2:  Shell heater temp (deg C) — 11.5

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4115 kg (9072 lb) as of 2/5/04  [SM(755) + FGB(2701) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:53am EST [= epoch]):

  • Mean altitude — 367.3 km
  • Apogee — 372.0 km
  • Perigee — 362.6 km
  • Period — 91.89 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.0007006
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29875

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.